OpenStudy (anonymous):

Solve –2(x – 1) = 14

OpenStudy (bahrom7893):

divide both sides by -2, then add 1 to both sides.

OpenStudy (anonymous):

would it be -8 or 8 as answer?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

No sorry it is either -8 or 8

OpenStudy (anonymous):

-2x+2=14 -2x=12 x=-6

OpenStudy (anonymous):

incorrect.

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

its -8 thanks